### v1

Abstract
Engagement in virtual learning is crucial for a variety of factors including learner satisfaction, performance, and compliance with learning programs, but measuring it is a challenging task. There is therefore considerable interest in utilizing artificial intelligence and affective computing to measure engagement in natural settings as well as on a large scale. This paper introduces a novel, privacy-preserving method for engagement measurement from videos. It uses facial landmarks, which carry no personally identifiable information, extracted from videos via the MediaPipe deep learning solution. The extracted facial landmarks are fed to a Spatial-Temporal Graph Convolutional Network (ST-GCN) to output the engagement level of the learner in the video. To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework based on transfer learning. Experimental results on two video student engagement measurement datasets show the superiority of the proposed method compared to previous methods with improved state-of-the-art on the EngageNet dataset with a %3.1 improvement in four-class engagement level classification accuracy and on the Online Student Engagement dataset with a %1.5 improvement in binary engagement classification accuracy. The relatively lightweight ST-GCN and its integration with the real-time MediaPipe deep learning solution make the proposed approach capable of being deployed on virtual learning platforms and measuring engagement in real time.
Keywords:
Engagement Measurement Graph Convolutional Network Ordinal Classification Transfer Learning.
1 Introduction
Engagement is intricately linked to the interaction of students’ attention and interest within a specific context [10243609]. It emerges not just as a consequence of pre-existing interests but also has the capability to cultivate new interests through consistent attention [hidi2006four]. This aspect is extremely crucial in educational environments, where the development and maintenance of engagement play a pivotal role in facilitating content comprehension [10243609]. Nonetheless, the challenges of measuring and sustaining engagement demand significant mental and emotional exertion from instructors. The progression in remote sensing technologies, Artificial Intelligence (AI), and affective computing holds promise for better measurement of engagement in diverse learning settings, including remote and virtual learning programs. Utilizing technologies such as facial expression recognition [liu2022graph], eye gaze tracking [wood2015rendering], and physiological stress detection [gedam2021review], instructors can more effectively monitor and enhance student engagement. This paper covers the development of AI algorithms for automated measurement of student engagement in virtual learning programs, taking into account the definition of engagement in educational psychology and its measurement methodologies.
Fredricks et al. [fredricks2004school] introduced three components of engagement: affective, behavioral, and cognitive engagement. Affective engagement relates to the emotions and attitudes a student has towards a task or its context, for instance, interest in a subject or feelings of excitement or confusion about a topic during a learning class [d2012dynamics]. Behavioral engagement involves the student’s active participation in a task, with indicators including looking at a computer screen (an on-task behavior) or playing with a mobile phone (an off-task behavior) during a learning class [ocumpaugh2015baker]. Cognitive engagement pertains to the student’s commitment to learning and their readiness to embrace challenges, encompassing cognitive outcomes such as recall, learning, and deep understanding [10243609].
Booth et al. [10243609] identified that the primary signals used to measure engagement components include facial, eye gaze, and body pose features, heart rate, Electro-Encephalo-Graphy, audio, and computer logs. Given that many of these signals can be captured by cameras and considering the widespread presence of cameras in personal computers and mobile phones—the devices commonly used for participation in virtual learning programs—video data has emerged as the predominant data modality for automatic AI-driven engagement measurement [karimah2022automatic, dewan2019engagement]. Consequently, most AI algorithms for engagement measurement have been designed and developed to analyze video data and infer the level of student engagement from the video [karimah2022automatic, dewan2019engagement, khan2022inconsistencies].
Engagement measurement methods are typically divided into end-to-end or feature-based categories [karimah2022automatic, dewan2019engagement, khan2022inconsistencies]. While feature-based approaches usually surpass end-to-end methods (that process raw videos) in performance [abedi2023affect, karimah2022automatic], they are constrained by the need to extract and identify the most effective features, often through trial and error. Typically, features from various modalities must be extracted from videos using multiple computer vision algorithms or deep neural networks. The extracted features are then analyzed by other deep neural networks or traditional machine learning methods to determine engagement [abedi2023affect, karimah2022automatic]. This requirement renders multi-modal feature-based approaches computationally heavy and challenging for practical implementation in real-world virtual learning platforms. These computationally intensive approaches render them impractical for local devices and necessitate the transfer of privacy-sensitive raw video from local devices to cloud servers to perform automated engagement measurement. Compared to raw videos, low-dimensional landmark information extracted from videos, such as facial landmarks, hand landmarks, and body joints [lugaresi2019mediapipe], offers not only more compact data but also encapsulates essential discriminative geometric features necessary for human affect [wei2023geometric] and behavior analysis [10144079], including engagement indicators. Most importantly, this geometric information is devoid of personally identifiable information [mishra2022skeletal].
In this paper, an alternative course away from conventional end-to-end and feature-based approaches is charted, and a novel engagement measurement technique using facial landmarks extracted from videos is presented. The proposed method is characterized by its privacy-preserving nature and computational efficiency. This work makes the following contributions:
  • This is the first instance in the field of video-based engagement measurement [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three, fwa2022fine] where facial landmarks extracted from videos, as the unique data modality, are analyzed using Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [yan2018spatial] to infer the engagement level in the video;
  • To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework utilizing transfer learning;
  • Extensive experiments conducted on two video-based engagement measurement datasets demonstrate the superiority of the proposed method over previous methods, achieving an improved state-of-the-art in engagement level classification accuracy.
The structure of this paper is organized as follows. Section 2 offers an overview of the relevant literature. This is succeeded by Section 3, which details the proposed methodology. Following this, Section 4 describes the experimental setup and discusses the results obtained with the proposed method. Lastly, Section 5 concludes the paper and proposes directions for future research.
2 Related Work
In line with the contributions of the proposed work in this paper, the literature review presented in this section serves a dual purpose: it discusses previous video-based engagement measurement techniques and explores the methods previously utilized variants of graph convolutional networks for the purpose of facial expression and affect analysis.
2.1 Engagement Measurement
Karimah et al. [karimah2022automatic] conducted a systematic review and explored the measurement of student engagement in virtual learning environments. They found that most engagement measurement methods primarily focused on the affective and behavioral components of engagement. These methods commonly employed datasets annotated by external observers [khan2022inconsistencies] to train feature-based and end-to-end machine-learning and deep-learning models. In the following, some of the relevant feature-based and end-to-end works on video-based engagement measurement are discussed.
In the domain of end-to-end engagement measurement techniques, deep neural networks analyze consecutive raw video frames to output the engagement level of the student in the video. These methods do not employ the extraction of handcrafted features from the videos; instead, the network is adept at autonomously learning to extract the most salient features directly from the videos, utilizing consecutive convolutional layers. The deep neural networks implemented in these end-to-end approaches include networks capable of video analysis, such as 3D Convolutional Neural Networks (3D CNNs), as well as combinations of 2D CNNs with sequential neural networks. For engagement measurement with 3D video input, networks such as C3D [gupta2016daisee, abedi2021improving], Inflated 3D CNN (I3D) [zhang2019novel], 3D DenseNet [mehta2022three], and Video Transformer [ai2022class] have been utilized. In subsequent methods, 2D CNNs and sequential models are jointly trained to automatically extract visual features from raw video frames, analyze the sequences of these features, and output the engagement level. Examples of this approach include the integration of 2D CNN with Long Short-Term Memory (LSTM) [gupta2016daisee], a combination of 2D Residual Neural Network (ResNet) and Temporal Convolutional Network (TCN) [abedi2021improving], ResNet paired with LSTM equipped with global attention [liao2021deep], and EfficientNet combined with LSTM and TCN [selim2022students].
The process of measuring engagement through feature-based techniques is divided into two stages. The first stage involves extracting behavioral and affective features from video frames, which depends on either specific domain knowledge or the use of pre-trained models for facial embedding extraction. Recent examples of pre-trained models used for extracting facial embedding features from videos for engagement measurement include Masked Autoencoder for facial video Representation LearnINg (MARLIN) [cai2023marlin] utilized by Singh et al. [engagenet] and Emotion Face Alignment Network (EmoFAN) [toisoul2021estimation] utilized by Abedi et al. [abedi2023affect]. OpenFace [baltrusaitis2018openface] is notably recognized for its effectiveness in extracting features associated with facial Action Units (AUs), eye movements, gaze direction, and the positioning of the head. OpenFace [baltrusaitis2018openface] has been widely used in feature-based engagement measurement techniques. The extracted features are then combined in various ways to create sequences of feature vectors for the video. These feature vector sequences, taken from consecutive video frames, form multivariate time series data. To analyze the multivariate time series, machine learning and deep learning models such as Bag-of-Words (BoW) models [abedi2023bag], variations of Recurrent Neural Networks (RNNs) [abedi2023detecting], TCNs [abedi2023affect], Transformers [engagenet], or ensemble models [tian2023predicting] are employed. These models are trained on the sequences of features to derive levels of engagement [copur2022engagement, whitehill2014faces, liao2021deep, niu2018automatic, thomas2018predicting, huang2019fine, fedotov2018multimodal, chen2019faceengage, booth2017toward, kaur2018prediction, wu2020advanced, ma2021automatic].
Prior studies have demonstrated the advantages of feature-based methods over end-to-end approaches in measuring engagement. Yet, feature-based methods are constrained by the need to extract and pinpoint the most discriminative features, often through a process of trial and error. Additionally, they necessitate the segmentation of videos into multiple segments (comprising multiple frames) and the extraction of diverse multi-modal features from all video segments using multiple deep neural networks. For example, techniques employing OpenFace 2.0 [baltrusaitis2018openface] for extracting features such as eye gaze, head pose, and Action Units (AUs) involve the use of multiple complex neural networks [wood2015rendering, zadeh2017convolutional, zhang2016joint]. This makes the methods time-intensive during the inference phase, as the features extracted must be further analyzed by another deep neural network to make a final inference regarding engagement [engagenet].
2.2 Graph-based Facial Affect and Expression Analysis
Liu et al. [liu2022graph] conducted a comprehensive review of the literature on graph-based methods for facial affect analysis. These methods typically take an image or a sequence of images as input and produce an affect classification or regression as output. Based on their review, Liu et al. [liu2022graph] proposed a pipeline for graph-based facial affect analysis, which includes (i) face preprocessing, (ii) graph-based affective representation, and (iii) graph-based relational reasoning. The preprocessing involves necessary steps such as face detection and registration. Graph-based affective representation involves defining the structure of the graph, i.e., nodes and edges. The graph structure can be spatial, or spatiotemporal when the input data is still images, or video frames, respectively. The graph structure can be at landmark-level, region-level, or AU-level with nodes representing facial landmarks, facial regions of interest, or facial AUs, respectively. In the relational reasoning step, the edges, nodes, their interrelations, and temporal dynamics are analyzed through relational reasoning machine-/deep-learning models to make an inference regarding affect. The machine-/deep-learning models used to analyze graph data include dynamic Bayesian networks, RNNs, CNNs, fully-connected neural networks, and non-temporal and temporal graph neural networks [liu2022graph].
The facial expression recognition method proposed by Zhou et al. [9191181] is among the methods where affective representation was performed by a spatiotemporal landmark-level and region-level graph. The method utilized thirty-four facial landmarks located around the eyes, lips, and cheeks, and the connections among these landmarks formed the intra-frame graph. These intra-frame connections were defined manually. To establish the inter-frame connections, each node in a frame was linked to its corresponding node in the subsequent frames. Two parallel ST-GCNs with analogous structures were trained: one on the x and y coordinates of the nodes, and the other on the histogram of orientation features derived from regions surrounding each node. The outputs from these two ST-GCNs were concatenated and then processed through a fully connected network for recognizing facial expressions in the input video. Among the disadvantages of this method are the independent training of the above-mentioned two ST-GCNs rather than joint learning and the manual definition of nodes and edges rather than allowing the ST-GCNs to learn the most relevant nodes and edges. Nodes were selected from Dlib’s facial landmarks [king_dlib] and additional nodes were created on the cheeks and between the eyes according to Dlib’s facial landmarks. The selection and creation of nodes were based on their relevance to certain AUs associated with the facial expressions that should be recognized.
Wei et al. [wei2023geometric] introduced a graph-based method for micro-expression recognition in video. The introduced approach included a dual-stream ST-GCN, focusing on the $x$ and $y$ coordinates of facial landmarks and the distance and angles between adjacent facial landmarks. Recognizing the association between AUs and micro-expressions, they integrated an AU-specific loss function into their neural network’s training process. Their methodology employed three different sets of facial landmarks as graph nodes, comprising sets with 14, 31, and Dlib’s [king_dlib] 68 facial landmarks. Notably, the set with 68 landmarks was the only one to include landmarks along the jawline. The experiments demonstrated the best results of micro-expression recognition when 14 facial landmark sets were used.
Leong et al. [fwa2022fine] introduced a method for detecting academic emotions through the use of facial landmarks and the roll, pitch, and yaw of head pose, utilizing spatial-temporal graph attention networks thereafter. Facial regions are cropped from video frames, and facial landmarks from these images are extracted, retaining those around the brow, eye, nose, and mouth, while excluding those that outline the face’s outer shape, with the reasoning being they lack correlation with affective states. These features are then inputted into spatial-temporal graph attention networks to determine academic emotions. A notable limitation of this approach is its reliance on multiple deep neural networks for extracting features, i.e., facial landmarks and head pose. The method achieved lower academic emotion detection accuracy when compared to prior feature-based methods.
2.3 Discussion
As reviewed in this section, while some methods have utilized facial landmarks and ST-GCNs for video-based facial affect and expression recognition, there has been no exploration in using facial landmarks or ST-GCNs specifically for video-based engagement measurement. It is important to note the fundamental differences between engagement and facial affect and expression making their measurement different. First, engagement is a multi-component state comprising behavioral and cognitive components in addition to its affective component. To illustrate, key indicators of behavioral engagement, such as being off-task or on-task, are determined by head pose, eye gaze, and blink rate. These indicators, and therefore engagement, cannot be effectively measured using methods designed solely for facial affect analysis. Second, engagement is not a constant state; it varies over time and should be measured at specific time resolutions where it remains stable and quantifiable. The ideal duration for measuring engagement is between 10 and 40 seconds, which is longer than the time resolution for facial affect analysis which is sometimes at frame level. Third, engagement measurement could involve recognizing an ordinal variable that indicates various levels of engagement, which differs from facial expression recognition, where the goal is to identify a categorical variable with no inherent order.
The limitations of existing engagement measurement approaches, discussed at the end of subsection 2.1, along with the key differences outlined above between facial affect analysis and engagement measurement, highlight a gap in the field of engagement measurement. To address this gap, we propose a simple yet effective engagement measurement technique. This method is also privacy-preserving and demonstrates improvements over existing methods on two video-based engagement measurement datasets.
3 Method
The input to the proposed method is a video sample of a student seated in front of a laptop or PC camera during a virtual learning session. The sequences of facial landmarks extracted from consecutive video frames through MediaPipe [lugaresi2019mediapipe, grishchenko2020attention] are analyzed by ST-GCN [yan2018spatial, 10144079] to output the engagement level of the student in the video.
3.1 Graph-based Representation
MediaPipe [lugaresi2019mediapipe], an open-source framework developed by Google, facilitates the construction of applied machine-learning pipelines for processing video data. The framework’s capabilities extend to a variety of tasks such as face tracking, hand and pose estimation, and facial landmark estimation, offering pre-trained models and tools suitable for real-time and streaming applications [latreche2023reliability]. Grishchenko et al. [grishchenko2020attention] introduced Attention Mesh, which was incorporated into MediaPipe [lugaresi2019mediapipe]. Attention Mesh is a pre-trained, end-to-end 3D facial landmark detection model that is both efficient and lightweight. It is enhanced with an attention mechanism focused on areas of semantic significance, such as the eyes, iris, and lips. Attention Mesh is capable of identifying 468 3D facial landmarks across the entire face and an additional 10 landmarks specifically for the iris. However, for the proposed method for engagement measurement, not all 478 landmarks are utilized. In line with existing research [liu2022graph], 68 3D facial landmarks that correspond to those identified by Dlib’s [king_dlib] framework, along with 10 3D landmarks from the iris, totaling 78 landmarks are utilized.
The 3D facial landmarks encapsulate crucial spatial-temporal information pertinent to head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation]—key features identified in prior engagement measurement studies [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three]. Consequently, there is no necessity for extracting additional handcrafted features from the video frames.
The $N$ 3D facial landmarks extracted from $T$ consecutive video frames are utilized to construct a spatiotemporal graph $G=(V,E)$. In this graph, the set of nodes $V=\{v_{ti}|t=1,\ldots,T,i=1,\ldots,N\}$ encompasses all the facial landmarks in a sequence. To construct $G$, first, the facial landmarks within one frame are connected with edges according to a connectivity structure based on Delaunay triangulation [liu2019facial] which is consistent with true facial muscle distribution and uniform for different subjects [liu2022graph]. Then each landmark will be connected to the same landmark in the consecutive frame.
3.2 Graph-based Reasoning
Based on the spatiotemporal graph of facial landmarks $G=(V,E)$ constructed above, an adjacency matrix $A$ is defined as an $N\times N$ matrix where the element at position $(i,j)$ is set to 1 if there is an edge connecting the $i^{th}$ and $j^{th}$ landmarks, and set to 0 otherwise. An identity matrix $I$, of the same dimensions as $A$, is created to represent self-connections. A spatial-temporal graph, as the basic element of an ST-GCN layer, is implemented as follows [kipf2016semi, yan2018spatial].
$f_{\text{out}}=\left(\Lambda^{-\frac{1}{2}}((A+I)\odot M)\Lambda^{-\frac{1}{2}}f_{\text{in}}W_{\text{spatial}}\right)W_{\text{temporal}}$ (1)
where $\Lambda^{ii}=\sum_{j}(A^{ij}+I^{ij})$. $M$ is a learnable weight matrix that enables scaling the contributions of a node’s feature to its neighboring nodes [yan2018spatial]. The input feature map, denoted $f_{\text{in}}$, is the raw coordinates of facial landmarks for the first layer of ST-GCN, and it represents the outputs from previous layers in subsequent layers of ST-GCN. The dimensionality of $f_{\text{in}}$ is $(C,N,T)$, where $C$ is the number of channels, for example, 3 in the initial input to the network corresponding to the $x$, $y$, and $z$ coordinates of the facial landmarks. In each layer of ST-GCN, initially, the spatial (intra-frame) convolution is applied to $f_{\text{in}}$ based on the weight matrix $W_{\text{spatial}}$, utilizing a standard 2D convolution with a kernel size of $1\times 1$. Subsequently, the resulting tensor is multiplied by the normalized adjacency matrix $\Lambda^{-\frac{1}{2}}((A+I)\odot M)\Lambda^{-\frac{1}{2}}$ across the spatial dimension. Afterward, the temporal (inter-frame) convolution, based on the weight matrix $W_{\text{temporal}}$, is applied to the tensor output from the spatial convolution. This convolution is a standard 2D convolution with a kernel size of $1\times\Gamma$, where $\Gamma$ signifies the temporal kernel size.
Following an adequate number of ST-GCN layers that perform spatial-temporal graph convolutions as outlined above, the resulting tensor undergoes 2D average pooling, and the final output of the network is generated by a final 2D convolution. This convolution employs a kernel size of $1\times 1$ and features an output channel dimensionality equal to the number of classes $K$, i.e., the number of engagement levels to be measured.
3.3 Ordinal Engagement classification through Transfer Learning
The model described above, with a final layer comprising $K$ output channels, tackles the engagement measurement problem as a categorical $K$-class classification problem without taking into account the ordinal nature of the engagement variable [yannakakis2018ordinal, whitehill2014faces, khan2022inconsistencies]. The model could harness the ordinality of the engagement variable to enhance its inferences. Drawing inspiration from [frank2001simple, abedi2023affect], a novel ordinal learning framework based on transfer learning is introduced as follows.
Training phase– The original $K$-level ordinal labels, $y=0,1,\ldots,K-1$, in the training set are converted into $K-1$ binary labels $y_{i}$ as follows: if $y>i$, then $y_{i}=1$; otherwise, $y_{i}=0$, for $i=0,1,\ldots,K-2$. Subsequently, $K-1$ binary classifiers are trained with the training set and the $K-1$ binary label sets described above. The training of binary classifiers is based on transfer learning, which proceeds as follows: Initially, a network is trained on the dataset with the original $K$-class labels, employing a regular final layer with $K$ output channels. After training, the ST-GCN layers of this network are frozen, and the final layer is removed. To this frozen network, $K-1$ separate untrained 2D convolution layers with a single output channel each are added, resulting in $K-1$ new networks. Each of these networks consists of a frozen sequence of ST-GCN layers followed by an untrained 2D convolution layer. These $K-1$ new networks are then trained on the entire dataset using the $K-1$ binary label sets described above. During this phase, only the final 2D convolution layers are subjected to training.
Inference phase– For the ordinal classification of a test sample, the sample is initially input into the pre-trained (and frozen) sequence of ST-GCN layers, followed by a 2D average pooling layer. The tensor obtained from this process is then input into $K-1$ pre-trained final 2D convolution layers, each yielding a probability estimate for the test sample being in the binary class $y_{t}=y_{i}$, where $i=0,1,\ldots,K-2$. Subsequently, these $K-1$ binary probability estimates are transformed into a single multi-class probability of the sample belonging to class $y=0,1,\ldots,K-1$, as follows [frank2001simple].
$p(y_{t}=k)=\begin{cases}1-p(y_{t}\geq 0),&\text{if }k=0,\\ p(y_{t}>k-1)-p(y_{t}\geq k),&\text{if }0<k<K-1,\\ p(y_{t}>K-2),&\text{if }k=K-1.\end{cases}$ (2)
Despite the increased training time for the ordinal model within the aforementioned ordinal learning framework, the final count of parameters in the ordinal model remains nearly identical to that of the original non-ordinal model.
4 Experiments
This section evaluates the performance of the proposed method relative to existing methods in video-based engagement measurement. It reports and discusses the results of multi-class and binary classification of engagement across two datasets.
4.1 Evaluation Metrics
Based on the engagement measurement problem at hand, several evaluation metrics are employed. In the context of multi-class engagement level classification, metrics such as accuracy and confusion matrix are reported. For binary engagement classification, accuracy, Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC), and Area Under the Curve of the Precision and Recall curve (AUC-PR) are utilized.
Table 1: The distribution of samples in not engaged, barely engaged, engaged, and highly engaged classes in the train and validation sets in the EngageNet dataset [engagenet].
Engagement Level Train Validation
Not engaged 1550 132
Barely engaged 1035 97
Engaged 1658 273
Highly Engaged 3740 569
Total 7983 1071
Table 2: The distribution of samples in engaged and not engaged classes in the train, validation, and test sets in the IIITB Online SE dataset [thomas2022automatic].
Engagement Level Train Validation Test
Not engaged 570 580 570
Engaged 2620 1080 720
Total 3190 1660 1290
4.2 Datasets
Experiments on two video-based engagement measurement datasets were conducted, each presenting unique challenges that further enabled the validation of the proposed method.
EngageNet The EngageNet dataset [engagenet], recognized as the largest publicly available dataset for video-based student engagement measurement, includes video recordings of 127 subjects participating in virtual learning courses. Each video sample has a duration of 10 seconds, with a frame rate of 30 fps and a resolution of $1280\times 720$ pixels. Three independent annotators annotated the subjects’ video recordings as four ordinal levels of engagement: not-engaged, barely engaged, engaged, and highly engaged. The inter-rater reliability among the annotators was high, with values of 0.73, 0.77, and 0.79, indicating a substantial agreement level. The dataset was divided into 7983, 1071, and 2257 samples for training, validation, and testing, respectively, using a subject-independent data split approach [engagenet]. However, only the training and validation sets were made available by the dataset creators and were utilized for the training and validation of predictive models in the experiments presented in this paper. Table 1 details the distribution of the four engagement classes within the training and validation sets.
Online SE: The Online SE dataset [thomas2022automatic], comprises videos of six students participating in online courses via the Zoom platform. These recordings span 10 seconds each, with a frame rate of 24 fps and a resolution of $220\times 155$ pixels. Two independent annotators annotated the video recordings as engaged and not engaged using the Human Expert Labeling Process (HELP) [aslan2017human, abedi2023bag]. The binary engagement labels achieved an inter-rater reliability score of 0.62, denoting a significant level of agreement. Following the exclusion of flawed videos [abedi2023bag], the dataset was segmented into 3190, 1660, and 1290 samples for training, validation, and testing, respectively. The distribution of engaged and not-engaged samples across the training, validation, and test sets is outlined in Table 2.
4.3 Experimental Setting
The sole information extracted from video frames is facial landmarks, which are analyzed by ST-GCN to determine the engagement level of the student in the video. Drawing inspiration from the pioneering works on body-joints-based action analysis [zheng2023skeleton, yan2018spatial], the proposed ST-GCN for facial-landmarks-based engagement measurement is structured as follows. The input facial landmarks are first processed through a batch normalization layer, followed by three consecutive ST-GCN layers with 64, 128, and 256 output channels, respectively. Residual connections are incorporated in the last two ST-GCN layers. A dropout rate of 0.1 is applied to each ST-GCN layer. The temporal kernel size in the ST-GCN layers is selected to be $9$. Subsequent to the ST-GCN layers, an average pooling layer is utilized, and its resulting tensor is directed into a 2D convolutional layer with 256 input channels and a number of output channels corresponding to the number of classes. A Softmax activation function then computes the probability estimates. In cases where engagement measurement is framed as a binary classification task, the terminal 2D convolution layer is configured with a single output channel, substituting Softmax with a Sigmoid function. The Sigmoid function is also employed for the individual binary classifiers within the ordinal learning framework detailed in subsection 3.3. The models are trained using the Adam optimizer with mini-batches of size $16$ and a learning rate of $0.001$ for $300$ epochs. The learning rate is decayed by a factor of $0.1$ every $100$ epochs.
Table 3: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet]: comparison of various feature sets and classification models from [engagenet] against two configurations of the proposed method - facial landmarks analyzed by ST-GCN and ordinal ST-GCN. Bolded values denote the best results.
Features Model Accuracy
Gaze LSTM 0.6125
Head Pose LSTM 0.6760
AU LSTM 0.6303
Gaze + Head Pose LSTM 0.6769
Gaze + Head Pose + AU LSTM 0.6704
Gaze CNN-LSTM 0.6060
Head Pose CNN-LSTM 0.6732
AU CNN-LSTM 0.6172
Gaze + AU CNN-LSTM 0.6275
Head Pose + AU CNN-LSTM 0.6751
Gaze + Head Pose + AU CNN-LSTM 0.6751
Gaze TCN 0.6256
Head Pose TCN 0.6611
AU TCN 0.6293
Gaze + Head Pose + AU TCN 0.6779
Gaze Transformer 0.5545
Gaze + Head Pose Transformer 0.6445
Gaze + Head Pose + AU Transformer 0.6910
Gaze + Head Pose + AU + MARLIN Transformer 0.6849
Facial Landmarks ST-GCN 0.6937
Facial Landmarks Ordinal ST-GCN 0.7124
the proposed method
4.4 Experimental Results
Table 3 presents the comparative results of two settings of the proposed method, a regular non-ordinal classifier and an ordinal classifier, with previous methods on the validation set of the EngageNet dataset [engagenet]. The engagement measurement in EngageNet [engagenet] is a four-class classification problem and the accuracy is reported as the evaluation metric. The previous methods used different combinations of OpenFace’s eye gaze, head pose, and AU features [baltrusaitis2018openface] along with MARLIN’s facial embedding features [cai2023marlin]. These features were classified by LSTM, CNN-LSTM, TCN, and Transformer, refer to [engagenet] for more details of the previous methods.
In the single feature configurations of previous methods in Table 3, head pose achieves better results compared to AUs, which is better than eye gaze. While head pose and eye gaze are indicators of behavioral engagement [abedi2023affect, ocumpaugh2015baker], AUs, which are associated with facial expressions and affect [toisoul2021estimation], are indicators of affective engagement. Combining these three features is always beneficial since engagement is a multi-component variable that can be measured by affective and behavioral indicators when the only available data is video [abedi2023affect]. Among classification models, utilizing more advanced models improves accuracy; the Transformer is better than TCN, which is better than CNN-LSTM and LSTM; however, it comes at the cost of increased computational complexity and the number of model parameters. It should also be noted that the feature extraction step for the three feature sets requires three complex deep learning models in OpenFace [baltrusaitis2018openface]. Considering only the classification models, the number of parameters in the Transformer, according to the architecture described in [engagenet], is 104,963,076, while the numbers of parameters in the proposed non-ordinal and ordinal ST-GCN are 861,688, and 861,431, respectively, indicating the efficiency of the proposed method. While being lightweight, the proposed method with ordinal STGCN improves the state-of-the-art by 3.1% compared to eye gaze, head pose, and AUs with Transformer.
Tables 4(a) and 4(b) depict the confusion matrices of the two configurations of the proposed method in Table 3, facial landmarks classified by non-ordinal and ordinal ST-GCN, respectively. As shown in Tables 4(a) and 4(b), incorporating ordinality significantly increases the number of correctly classified samples in the first three classes and results in a 2.7% improvement in accuracy compared to its non-ordinal counterpart.
Table 4: Confusion matrices of the proposed method using facial landmarks with (a) non-ordinal and (b) ordinal ST-GCN on the validation set of the EngageNet dataset [engagenet].
Class 1 2 3 4
1 99 14 13 6
2 10 29 33 25
3 9 19 100 145
4 5 4 45 515
Class 1 2 3 4
1 104 12 10 6
2 15 36 24 22
3 10 26 112 125
4 9 5 44 511
Table 5 displays the results of different variants of the proposed method on the validation set of the EngageNet [engagenet] dataset. In the first two variants, the $x$, $y$, and $z$ coordinates of facial landmarks are converted into multivariate time series and analyzed by an LSTM and TCN. The LSTM includes four unidirectional layers with 256 neurons in hidden units and is followed by a $256\times 4$ fully connected layer. The parameters of the TCN are as follows: the number of layers, number of filters, kernel size, and dropout rate are 8, 64, 8, and 0.05, respectively. While their results are acceptable and better than most of the earlier methods in Table 3, they cannot outperform ST-GCN, the last two columns of Table 3. The fact that the accuracy of the LSTM and TCN in Table 5 is higher than those in Table 3 signifies the efficiency of facial landmarks for engagement measurement. In the third variant, the $z$ coordinates of facial landmarks are disregarded, and the decrease in accuracy indicates the importance of the $z$ coordinates for engagement measurement. Temporal kernel sizes other than 9 in the fourth and fifth variants have a negative impact on the results. When engagement measurement is performed using every 2, 4, 8, and 16 frames, instead of every frame, there is a slight decrease in accuracy. However, this is a trade-off between accuracy and computation since reducing the frame rate corresponds to a reduction in computation.
Table 5: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet] for different variants of the proposed method.
Variant Accuracy
ST-GCN is replaced with LSTM 0.6847
ST-GCN is replaced with TCN 0.6813
Only $x$ and $y$ coordinates of joints are used 0.6655
A temporal kernel of 3 is used 0.6748
A temporal kernel of 15 is used 0.6841
Every 2 frames are used 0.6813
Every 4 frames are used 0.6907
Every 8 frames are used 0.6907
Every 16 frames are used 0.6841
Table 6 presents the results of the proposed method in comparison to previous methods on the test set of the Online SE dataset [thomas2022automatic]. The earlier methods listed in Table 6 are feature-based, extracting affective and behavioral features from video frames and performing binary classification of the features using TCN in [thomas2018predicting], TCN in [abedi2023affect], LSTM with attention in [chen2019faceengage], and BoW in [abedi2023bag]. Given that engagement measurement in the Online SE dataset [thomas2022automatic] is posed as a binary classification problem, implementing the ordinal version of the proposed method is not required. The proposed method, employing facial landmarks with ST-GCN, attains the highest accuracy and AUC PR.
Table 6: Binary classification accuracy, AUC ROC, and AUC PR of engagement on the test set of the Online SE dataset [thomas2022automatic]: comparison of previous methods against the proposed method. Bolded values denote the best results.
Method Accuracy AUC ROC AUC PR
[thomas2018predicting] 0.7803 0.8764 0.8008
[abedi2023affect] 0.7637 0.8710 0.7980
[chen2019faceengage] 0.7475 0.8890 0.7973
[abedi2023bag] 0.8191 0.8926 0.9018
Proposed 0.8315 0.8806 0.9131
5 Conclusion and Future Works
Our research led to the development of a novel deep-learning framework for student engagement measurement. In the proposed framework, sequences of facial landmarks are extracted from consecutive video frames and analyzed by ordinal ST-GCNs to make inferences regarding the engagement level of the subject in the video. The successful application of our model to the EngageNet [engagenet] and Online SE [thomas2022automatic] datasets not only confirms its efficacy but also establishes a new standard in engagement level classification accuracy, outperforming existing methods. As the sole input information to the developed engagement measurement models, the 3D facial landmarks contain information about head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation], which are the key indicators of behavioral and affective engagement. The relatively lightweight ST-GCN and its integration with real-time MediaPipe make the proposed framework capable of being deployed on virtual learning platforms and measuring engagement in real time. Despite these advancements, the proposed method is not without limitations. The interpretability and explainability of the proposed method could be improved through techniques such as Gradient-based Class Activation Maps [zheng2023skeleton] for ST-GCNs [9746621]. This would facilitate an understanding of which facial landmarks and specific timestamps contribute to certain levels of engagement. To improve the performance of the proposed method, the following direction could be investigated: analyzing facial landmarks with more advanced ST-GCN, which are equipped with attention mechanism [lin2023actionlet], applying augmentation techniques to video data before facial landmark extraction [abedi2023cross] or to facial landmark data to improve the generalizability of the ST-GCNs, and incorporating hand landmarks and body joints extracted from video into the developed facial-landmarks-based framework.
Acknowledgment The authors express their sincere thanks to the Multimodal Perception Lab at the International Institute of Information Technology, Bangalore, India, for their generosity in providing the Online SE dataset, which was instrumental in the execution of our experiments.
[…]

### v2

@@
-
+Engagement Measurement Based on Facial Landmarks and Spatial-Temporal Graph Convolutional Networks
 Abstract
-Engagement in virtual learning is crucial for a variety of factors including learner satisfaction, performance, and compliance with learning programs, but measuring it is a challenging task.
+Engagement in virtual learning is crucial for a variety of factors including student satisfaction, performance, and compliance with learning programs, but measuring it is a challenging task.
 There is therefore considerable interest in utilizing artificial intelligence and affective computing to measure engagement in natural settings as well as on a large scale.
@@
 It uses facial landmarks, which carry no personally identifiable information, extracted from videos via the MediaPipe deep learning solution.
-The extracted facial landmarks are fed to a Spatial-Temporal Graph Convolutional Network (ST-GCN) to output the engagement level of the learner in the video.
+The extracted facial landmarks are fed to Spatial-Temporal Graph Convolutional Networks (ST-GCNs) to output the engagement level of the student in the video.
 To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework based on transfer learning.
-Experimental results on two video student engagement measurement datasets show the superiority of the proposed method compared to previous methods with improved state-of-the-art on the EngageNet dataset with a %3.1 improvement in four-class engagement level classification accuracy and on the Online Student Engagement dataset with a %1.5 improvement in binary engagement classification accuracy.
-The relatively lightweight ST-GCN and its integration with the real-time MediaPipe deep learning solution make the proposed approach capable of being deployed on virtual learning platforms and measuring engagement in real time.
+Experimental results on two video student engagement measurement datasets show the superiority of the proposed method compared to previous methods with improved state-of-the-art on the EngageNet dataset with a 3.1% improvement in four-class engagement level classification accuracy and on the Online Student Engagement dataset with a 1.5% improvement in binary engagement classification accuracy.
+Gradient-weighted Class Activation Mapping (Grad-CAM) was applied to the developed ST-GCNs to interpret the engagement measurements obtained by the proposed method in both the spatial and temporal domains.
+The relatively lightweight and fast ST-GCN and its integration with the real-time MediaPipe make the proposed approach capable of being deployed on virtual learning platforms and measuring engagement in real-time.
 Keywords:
@@
 1 Introduction
-Engagement is intricately linked to the interaction of students’ attention and interest within a specific context [10243609].
-It emerges not just as a consequence of pre-existing interests but also has the capability to cultivate new interests through consistent attention [hidi2006four].
-This aspect is extremely crucial in educational environments, where the development and maintenance of engagement play a pivotal role in facilitating content comprehension [10243609].
-Nonetheless, the challenges of measuring and sustaining engagement demand significant mental and emotional exertion from instructors.
-The progression in remote sensing technologies, Artificial Intelligence (AI), and affective computing holds promise for better measurement of engagement in diverse learning settings, including remote and virtual learning programs.
-Utilizing technologies such as facial expression recognition [liu2022graph], eye gaze tracking [wood2015rendering], and physiological stress detection [gedam2021review], instructors can more effectively monitor and enhance student engagement.
-This paper covers the development of AI algorithms for automated measurement of student engagement in virtual learning programs, taking into account the definition of engagement in educational psychology and its measurement methodologies.
+Engagement is key in education, blending students’ attention and interest within a learning context [1].
+It not only stems from existing interests but also fosters new ones through sustained attention [2], essential for content comprehension and engagement development [1].
+However, measuring and upholding engagement is challenging, and demands significant effort from educators.
+The advent of remote sensing, Artificial Intelligence (AI), and affective computing offers new avenues for accurately measuring engagement across various learning environments, including virtual learning platforms.
+Technologies such as facial expression recognition [3] and eye gaze tracking [4] enable more precise monitoring and enhancement of student engagement.
+This paper explores the development of AI algorithms for automated engagement measurement in virtual learning, taking into account the definition of engagement in educational psychology and its measurement methodologies.
 Fredricks et al.
-[fredricks2004school] introduced three components of engagement: affective, behavioral, and cognitive engagement.
-Affective engagement relates to the emotions and attitudes a student has towards a task or its context, for instance, interest in a subject or feelings of excitement or confusion about a topic during a learning class [d2012dynamics].
-Behavioral engagement involves the student’s active participation in a task, with indicators including looking at a computer screen (an on-task behavior) or playing with a mobile phone (an off-task behavior) during a learning class [ocumpaugh2015baker].
-Cognitive engagement pertains to the student’s commitment to learning and their readiness to embrace challenges, encompassing cognitive outcomes such as recall, learning, and deep understanding [10243609].
+[5] defined engagement through affective, behavioral, and cognitive components.
+Affective engagement relates to students’ emotions and attitudes towards their tasks, including interest or feelings during a learning class [6].
+Behavioral engagement refers to active participation, such as focusing on a computer screen or not being distracted by a phone [7].
+Cognitive engagement deals with a student’s dedication to learning and tackling challenges, affecting outcomes such as recall and understanding [1].
 Booth et al.
-[10243609] identified that the primary signals used to measure engagement components include facial, eye gaze, and body pose features, heart rate, Electro-Encephalo-Graphy, audio, and computer logs.
-Given that many of these signals can be captured by cameras and considering the widespread presence of cameras in personal computers and mobile phones—the devices commonly used for participation in virtual learning programs—video data has emerged as the predominant data modality for automatic AI-driven engagement measurement [karimah2022automatic, dewan2019engagement].
-Consequently, most AI algorithms for engagement measurement have been designed and developed to analyze video data and infer the level of student engagement from the video [karimah2022automatic, dewan2019engagement, khan2022inconsistencies].
-Engagement measurement methods are typically divided into end-to-end or feature-based categories [karimah2022automatic, dewan2019engagement, khan2022inconsistencies].
-While feature-based approaches usually surpass end-to-end methods (that process raw videos) in performance [abedi2023affect, karimah2022automatic], they are constrained by the need to extract and identify the most effective features, often through trial and error.
-Typically, features from various modalities must be extracted from videos using multiple computer vision algorithms or deep neural networks.
-The extracted features are then analyzed by other deep neural networks or traditional machine learning methods to determine engagement [abedi2023affect, karimah2022automatic].
-This requirement renders multi-modal feature-based approaches computationally heavy and challenging for practical implementation in real-world virtual learning platforms.
-These computationally intensive approaches render them impractical for local devices and necessitate the transfer of privacy-sensitive raw video from local devices to cloud servers to perform automated engagement measurement.
-Compared to raw videos, low-dimensional landmark information extracted from videos, such as facial landmarks, hand landmarks, and body joints [lugaresi2019mediapipe], offers not only more compact data but also encapsulates essential discriminative geometric features necessary for human affect [wei2023geometric] and behavior analysis [10144079], including engagement indicators.
-Most importantly, this geometric information is devoid of personally identifiable information [mishra2022skeletal].
+[1] identified that engagement is measured through various signals such as facial expressions, eye movements, posture, heart rate, brain activity, audio, and digital interactions.
+Cameras, prevalent in devices for online learning, make video a key data modality for AI-based engagement measurement [8, 9].
+Therefore, AI techniques predominantly focus on video to measure student engagement [8, 9, 10].
+AI-driven engagement measurement methods are divided into end-to-end and feature-based approaches.
+While feature-based methods generally outperform end-to-end approaches, which process raw videos, they require extensive identification of effective features through trial and error [8, 9, 10].
+Extracting multi-modal features from videos with multiple computer vision algorithms or neural networks [11, 12], and their subsequent analysis by deep neural networks to measure engagement, render these methods computationally demanding [13, 8].
+Such computational requirements limit their use on local devices and necessitate the transfer of privacy-sensitive video data to cloud servers for engagement measurement.
+In contrast, extracting low-dimensional landmark information, such as facial and hand landmarks, not only provides more compact data but also captures essential geometric features for affect and behavior analysis, including engagement, without personal identifiers [14, 15, 16].
+Previous feature-based engagement measurement approaches extracted features such as head pose, facial Action Units (AUs), iris and gaze features, and affect and emotion features, which are indicators of the behavioral and affective components of engagement [8, 9, 10].
+The literature has demonstrated the success of facial landmarks in capturing these aforementioned features [17, 18, 19, 20, 3].
+A model such as Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [21], capable of analyzing spatial-temporal facial landmarks and learning these aforementioned features inherently, can reduce the need for raw facial videos.
+An approach based on facial landmarks prioritizes privacy and reduces computational demands, making it more practical for real-time engagement measurement.
 In this paper, an alternative course away from conventional end-to-end and feature-based approaches is charted, and a novel engagement measurement technique using facial landmarks extracted from videos is presented.
@@
 This work makes the following contributions:
-  • This is the first instance in the field of video-based engagement measurement [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three, fwa2022fine] where facial landmarks extracted from videos, as the unique data modality, are analyzed using Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [yan2018spatial] to infer the engagement level in the video;
+  • This marks the first instance in video-based engagement measurement [8, 9, 22, 23, 13, 24, 25] where facial landmarks, as the single data modality extracted from videos, are analyzed through ST-GCNs [21] to infer the engagement level in the video;
   • To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework utilizing transfer learning;
   • Extensive experiments conducted on two video-based engagement measurement datasets demonstrate the superiority of the proposed method over previous methods, achieving an improved state-of-the-art in engagement level classification accuracy.
-The structure of this paper is organized as follows.
-Section 2 offers an overview of the relevant literature.
-This is succeeded by Section 3, which details the proposed methodology.
-Following this, Section 4 describes the experimental setup and discusses the results obtained with the proposed method.
-Lastly, Section 5 concludes the paper and proposes directions for future research.
+For explainability, Gradient-weighted Class Activation Mapping (Grad-CAM) is applied to the developed ST-GCNs to interpret the engagement measurements obtained by the proposed method in both spatial and temporal domains.
 2 Related Work
-In line with the contributions of the proposed work in this paper, the literature review presented in this section serves a dual purpose: it discusses previous video-based engagement measurement techniques and explores the methods previously utilized variants of graph convolutional networks for the purpose of facial expression and affect analysis.
+The literature review in this paper serves two main purposes: discussing past video-based engagement measurement techniques and examining the use of graph convolutional networks for facial expression and affect analysis.
 2.1 Engagement Measurement
 Karimah et al.
-[karimah2022automatic] conducted a systematic review and explored the measurement of student engagement in virtual learning environments.
-They found that most engagement measurement methods primarily focused on the affective and behavioral components of engagement.
-These methods commonly employed datasets annotated by external observers [khan2022inconsistencies] to train feature-based and end-to-end machine-learning and deep-learning models.
+[8] conducted a systematic review on measuring student engagement in virtual learning environments, revealing a focus on affective and behavioral components of engagement.
+Most methods utilize datasets annotated by external observers [10] to train both feature-based and end-to-end models.
 In the following, some of the relevant feature-based and end-to-end works on video-based engagement measurement are discussed.
 In the domain of end-to-end engagement measurement techniques, deep neural networks analyze consecutive raw video frames to output the engagement level of the student in the video.
-These methods do not employ the extraction of handcrafted features from the videos; instead, the network is adept at autonomously learning to extract the most salient features directly from the videos, utilizing consecutive convolutional layers.
-The deep neural networks implemented in these end-to-end approaches include networks capable of video analysis, such as 3D Convolutional Neural Networks (3D CNNs), as well as combinations of 2D CNNs with sequential neural networks.
-For engagement measurement with 3D video input, networks such as C3D [gupta2016daisee, abedi2021improving], Inflated 3D CNN (I3D) [zhang2019novel], 3D DenseNet [mehta2022three], and Video Transformer [ai2022class] have been utilized.
-In subsequent methods, 2D CNNs and sequential models are jointly trained to automatically extract visual features from raw video frames, analyze the sequences of these features, and output the engagement level.
-Examples of this approach include the integration of 2D CNN with Long Short-Term Memory (LSTM) [gupta2016daisee], a combination of 2D Residual Neural Network (ResNet) and Temporal Convolutional Network (TCN) [abedi2021improving], ResNet paired with LSTM equipped with global attention [liao2021deep], and EfficientNet combined with LSTM and TCN [selim2022students].
-The process of measuring engagement through feature-based techniques is divided into two stages.
-The first stage involves extracting behavioral and affective features from video frames, which depends on either specific domain knowledge or the use of pre-trained models for facial embedding extraction.
-Recent examples of pre-trained models used for extracting facial embedding features from videos for engagement measurement include Masked Autoencoder for facial video Representation LearnINg (MARLIN) [cai2023marlin] utilized by Singh et al.
-[engagenet] and Emotion Face Alignment Network (EmoFAN) [toisoul2021estimation] utilized by Abedi et al.
-[abedi2023affect].
-OpenFace [baltrusaitis2018openface] is notably recognized for its effectiveness in extracting features associated with facial Action Units (AUs), eye movements, gaze direction, and the positioning of the head.
-OpenFace [baltrusaitis2018openface] has been widely used in feature-based engagement measurement techniques.
-The extracted features are then combined in various ways to create sequences of feature vectors for the video.
-These feature vector sequences, taken from consecutive video frames, form multivariate time series data.
-To analyze the multivariate time series, machine learning and deep learning models such as Bag-of-Words (BoW) models [abedi2023bag], variations of Recurrent Neural Networks (RNNs) [abedi2023detecting], TCNs [abedi2023affect], Transformers [engagenet], or ensemble models [tian2023predicting] are employed.
-These models are trained on the sequences of features to derive levels of engagement [copur2022engagement, whitehill2014faces, liao2021deep, niu2018automatic, thomas2018predicting, huang2019fine, fedotov2018multimodal, chen2019faceengage, booth2017toward, kaur2018prediction, wu2020advanced, ma2021automatic].
-Prior studies have demonstrated the advantages of feature-based methods over end-to-end approaches in measuring engagement.
-Yet, feature-based methods are constrained by the need to extract and pinpoint the most discriminative features, often through a process of trial and error.
-Additionally, they necessitate the segmentation of videos into multiple segments (comprising multiple frames) and the extraction of diverse multi-modal features from all video segments using multiple deep neural networks.
-For example, techniques employing OpenFace 2.0 [baltrusaitis2018openface] for extracting features such as eye gaze, head pose, and Action Units (AUs) involve the use of multiple complex neural networks [wood2015rendering, zadeh2017convolutional, zhang2016joint].
-This makes the methods time-intensive during the inference phase, as the features extracted must be further analyzed by another deep neural network to make a final inference regarding engagement [engagenet].
+These methods do not employ the extraction of handcrafted features from the videos; instead, the network is adept at autonomously learning to extract the most useful features directly from the videos, utilizing consecutive convolutional layers.
+The deep neural networks implemented in these end-to-end approaches include networks capable of video analysis, such as 3D Convolutional Neural Networks (3D CNNs) and Video Transformers [26, 27, 28], as well as combinations of 2D CNNs with sequential neural networks such as Long Short-Term Memory (LSTM) and Temporal Convolutional Network (TCN) [26, 27, 29, 30].
+The process of measuring engagement through feature-based techniques involves two stages.
+Initially, behavioral and affective features are extracted from video frames, relying on either domain-specific knowledge or pre-trained models for facial embedding extraction.
+OpenFace [31] is notably effective for its comprehensive feature extraction capabilities, including AUs, eye movements, gaze direction, and head positioning, and is widely applied in engagement measurement [13, 32, 33, 12].
+Examples of facial embedding include the Masked Autoencoder for facial video Representation LearnINg (MARLIN) [11], utilized by Singh et al.
+[12], and the Emotion Face Alignment Network (EmoFAN) [19], used by Abedi et al.
+[13].
+Subsequently, to analyze these extracted features and infer engagement, various machine learning and deep learning models are employed, such as Bag-of-Words (BoW) [32], Recurrent Neural Network (RNN) variations [24], Temporal Convolutional Networks (TCNs) [13, 33], Transformers [12, 34], and ensemble models [35].
 2.2 Graph-based Facial Affect and Expression Analysis
 Liu et al.
-[liu2022graph] conducted a comprehensive review of the literature on graph-based methods for facial affect analysis.
+[3] conducted a comprehensive review of the literature on graph-based methods for facial affect analysis.
 These methods typically take an image or a sequence of images as input and produce an affect classification or regression as output.
 Based on their review, Liu et al.
-[liu2022graph] proposed a pipeline for graph-based facial affect analysis, which includes (i) face preprocessing, (ii) graph-based affective representation, and (iii) graph-based relational reasoning.
-The preprocessing involves necessary steps such as face detection and registration.
+[3] proposed a pipeline for graph-based facial affect analysis, which includes (i) face preprocessing, (ii) graph-based affective representation, and (iii) graph-based relational reasoning.
+Preprocessing involves steps such as face detection and registration.
 Graph-based affective representation involves defining the structure of the graph, i.e., nodes and edges.
-The graph structure can be spatial, or spatiotemporal when the input data is still images, or video frames, respectively.
-The graph structure can be at landmark-level, region-level, or AU-level with nodes representing facial landmarks, facial regions of interest, or facial AUs, respectively.
-In the relational reasoning step, the edges, nodes, their interrelations, and temporal dynamics are analyzed through relational reasoning machine-/deep-learning models to make an inference regarding affect.
-The machine-/deep-learning models used to analyze graph data include dynamic Bayesian networks, RNNs, CNNs, fully-connected neural networks, and non-temporal and temporal graph neural networks [liu2022graph].
-The facial expression recognition method proposed by Zhou et al.
-[9191181] is among the methods where affective representation was performed by a spatiotemporal landmark-level and region-level graph.
-The method utilized thirty-four facial landmarks located around the eyes, lips, and cheeks, and the connections among these landmarks formed the intra-frame graph.
-These intra-frame connections were defined manually.
-To establish the inter-frame connections, each node in a frame was linked to its corresponding node in the subsequent frames.
-Two parallel ST-GCNs with analogous structures were trained: one on the x and y coordinates of the nodes, and the other on the histogram of orientation features derived from regions surrounding each node.
-The outputs from these two ST-GCNs were concatenated and then processed through a fully connected network for recognizing facial expressions in the input video.
-Among the disadvantages of this method are the independent training of the above-mentioned two ST-GCNs rather than joint learning and the manual definition of nodes and edges rather than allowing the ST-GCNs to learn the most relevant nodes and edges.
-Nodes were selected from Dlib’s facial landmarks [king_dlib] and additional nodes were created on the cheeks and between the eyes according to Dlib’s facial landmarks.
-The selection and creation of nodes were based on their relevance to certain AUs associated with the facial expressions that should be recognized.
+The graph structure can be spatial or spatiotemporal depending on whether the input data is still images or videos.
+The graph structure can be at the landmark level, region level, or AU level, with nodes representing facial landmarks, facial regions of interest, or facial AUs, respectively.
+In the relational reasoning step, the edges, nodes, their interrelations, and temporal dynamics are analyzed through relational reasoning machine-learning or deep-learning models to make inferences regarding affect.
+Models used to analyze graph data include dynamic Bayesian networks, RNNs, CNNs, fully-connected neural networks, and non-temporal and temporal graph neural networks [3].
+Zhou et al.
+[36] proposed a facial expression recognition method based on spatiotemporal landmark-level and region-level graphs.
+The intra-frame graph is formed by the connections among thirty-four facial landmarks located around the eyes, lips, and cheeks.
+The definition of these intra-frame connections was done manually.
+Inter-frame connections were established by linking each node in one frame to its corresponding node in the following frame.
+Two parallel ST-GCNs with analogous structures were trained; one on the nodes’ $x$- and $y$-coordinates, and another on their histogram of orientation features.
+ST-GCNs’ outputs were concatenated and processed by a fully connected network to identify facial expressions.
+This method’s disadvantages include independent training of the two ST-GCNs rather than joint learning, and manual definition of nodes and edges.
 Wei et al.
-[wei2023geometric] introduced a graph-based method for micro-expression recognition in video.
-The introduced approach included a dual-stream ST-GCN, focusing on the $x$ and $y$ coordinates of facial landmarks and the distance and angles between adjacent facial landmarks.
-Recognizing the association between AUs and micro-expressions, they integrated an AU-specific loss function into their neural network’s training process.
-Their methodology employed three different sets of facial landmarks as graph nodes, comprising sets with 14, 31, and Dlib’s [king_dlib] 68 facial landmarks.
+[15] presented a graph-based method for micro-expression recognition in video.
+The method included a dual-stream ST-GCN, focusing on the $x$ and $y$ coordinates of facial landmarks and the distance and angles between adjacent facial landmarks.
+An AU-specific loss function was incorporated into the neural network’s training process in order to incorporate the association between AUs and micro-expressions.
+Their methodology employed three different sets of facial landmarks as graph nodes, comprising sets with 14, 31, and Dlib’s [37] 68 facial landmarks.
 Notably, the set with 68 landmarks was the only one to include landmarks along the jawline.
-The experiments demonstrated the best results of micro-expression recognition when 14 facial landmark sets were used.
+The experiments demonstrated the best results for micro-expression recognition when 14 facial landmark sets were used.
 Leong et al.
-[fwa2022fine] introduced a method for detecting academic emotions through the use of facial landmarks and the roll, pitch, and yaw of head pose, utilizing spatial-temporal graph attention networks thereafter.
-Facial regions are cropped from video frames, and facial landmarks from these images are extracted, retaining those around the brow, eye, nose, and mouth, while excluding those that outline the face’s outer shape, with the reasoning being they lack correlation with affective states.
-These features are then inputted into spatial-temporal graph attention networks to determine academic emotions.
+[25] introduced a method employing spatial-temporal graph attention networks to analyze facial landmarks and head poses, aimed at identifying academic emotions.
+The facial landmarks used were those around the eyebrow, eye, nose, and mouth excluding those that outline the face’s outer shape, with the reasoning being they lack correlation with affective states.
 A notable limitation of this approach is its reliance on multiple deep neural networks for extracting features, i.e., facial landmarks and head pose.
-The method achieved lower academic emotion detection accuracy when compared to prior feature-based methods.
+The method achieved lower academic emotion detection accuracy when compared to prior feature-based methods [13].
 2.3 Discussion
-As reviewed in this section, while some methods have utilized facial landmarks and ST-GCNs for video-based facial affect and expression recognition, there has been no exploration in using facial landmarks or ST-GCNs specifically for video-based engagement measurement.
-It is important to note the fundamental differences between engagement and facial affect and expression making their measurement different.
-First, engagement is a multi-component state comprising behavioral and cognitive components in addition to its affective component.
+As reviewed in this section, while some methods have used facial landmarks and ST-GCNs to recognize facial affect and expression in videos, there has been no exploration of the use of these methods to measure engagement.
+The fundamental differences between engagement and facial affect and expression make their measurement different.
+First, engagement is a multi-component state comprising behavioral, affective, and cognitive components.
 To illustrate, key indicators of behavioral engagement, such as being off-task or on-task, are determined by head pose, eye gaze, and blink rate.
@@
 Second, engagement is not a constant state; it varies over time and should be measured at specific time resolutions where it remains stable and quantifiable.
-The ideal duration for measuring engagement is between 10 and 40 seconds, which is longer than the time resolution for facial affect analysis which is sometimes at frame level.
-Third, engagement measurement could involve recognizing an ordinal variable that indicates various levels of engagement, which differs from facial expression recognition, where the goal is to identify a categorical variable with no inherent order.
-The limitations of existing engagement measurement approaches, discussed at the end of subsection 2.1, along with the key differences outlined above between facial affect analysis and engagement measurement, highlight a gap in the field of engagement measurement.
-To address this gap, we propose a simple yet effective engagement measurement technique.
-This method is also privacy-preserving and demonstrates improvements over existing methods on two video-based engagement measurement datasets.
+An ideal measurement duration for engagement is between ten and forty seconds, which is longer than the time resolution for facial affect analysis, which sometimes occurs at the frame level.
+Third, engagement measurement could involve recognizing an ordinal variable that indicates levels of engagement, as opposed to facial expression recognition, which identifies categorical variables without inherent order.
+Existing engagement measurement approaches face limitations due to the necessity of employing multiple deep neural networks for the extraction and analysis of multi-modal features.
+Coupled with the significant differences between facial affect analysis and engagement measurement as outlined above, these limitations underscore a gap in the field.
+In response, we introduce a straightforward yet effective method for engagement measurement.
+This method is lightweight and fast, preserving privacy while also demonstrating improvements over current methodologies across two video-based engagement measurement datasets.
 3 Method
 The input to the proposed method is a video sample of a student seated in front of a laptop or PC camera during a virtual learning session.
-The sequences of facial landmarks extracted from consecutive video frames through MediaPipe [lugaresi2019mediapipe, grishchenko2020attention] are analyzed by ST-GCN [yan2018spatial, 10144079] to output the engagement level of the student in the video.
+The sequences of facial landmarks extracted from consecutive video frames through MediaPipe [14, 20] are analyzed by ST-GCN [21, 16] to output the engagement level of the student in the video.
 3.1 Graph-based Representation
-MediaPipe [lugaresi2019mediapipe], an open-source framework developed by Google, facilitates the construction of applied machine-learning pipelines for processing video data.
-The framework’s capabilities extend to a variety of tasks such as face tracking, hand and pose estimation, and facial landmark estimation, offering pre-trained models and tools suitable for real-time and streaming applications [latreche2023reliability].
-Grishchenko et al.
-[grishchenko2020attention] introduced Attention Mesh, which was incorporated into MediaPipe [lugaresi2019mediapipe].
-Attention Mesh is a pre-trained, end-to-end 3D facial landmark detection model that is both efficient and lightweight.
-It is enhanced with an attention mechanism focused on areas of semantic significance, such as the eyes, iris, and lips.
-Attention Mesh is capable of identifying 468 3D facial landmarks across the entire face and an additional 10 landmarks specifically for the iris.
-However, for the proposed method for engagement measurement, not all 478 landmarks are utilized.
-In line with existing research [liu2022graph], 68 3D facial landmarks that correspond to those identified by Dlib’s [king_dlib] framework, along with 10 3D landmarks from the iris, totaling 78 landmarks are utilized.
-The 3D facial landmarks encapsulate crucial spatial-temporal information pertinent to head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation]—key features identified in prior engagement measurement studies [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three].
+The MediaPipe deep learning solution [14], a framework that is both real-time and cross-platform, is employed for the extraction of facial landmarks from video.
+Incorporated within MediaPipe, Attention Mesh [20] is adept at detecting 468 3D facial landmarks throughout the face and an extra 10 landmarks for the iris.
+However, not all 478 landmarks are employed in the proposed engagement measurement method.
+Consistent with existing studies [3], only 68 of the 3D facial landmarks, which match those identified by the Dlib framework [37], in addition to the 10 3D iris landmarks, making a total of 78 landmarks, are utilized.
+The 3D facial landmarks encapsulate crucial spatial-temporal information pertinent to head pose [17], AUs [18], eye gaze [20], and affect [3, 19]—key features identified in prior engagement measurement studies [8, 9, 10, 22, 23, 13, 24, 33, 26, 27].
 Consequently, there is no necessity for extracting additional handcrafted features from the video frames.
@@
 In this graph, the set of nodes $V=\{v_{ti}|t=1,\ldots,T,i=1,\ldots,N\}$ encompasses all the facial landmarks in a sequence.
-To construct $G$, first, the facial landmarks within one frame are connected with edges according to a connectivity structure based on Delaunay triangulation [liu2019facial] which is consistent with true facial muscle distribution and uniform for different subjects [liu2022graph].
+To construct $G$, first, the facial landmarks within one frame are connected with edges according to a connectivity structure based on Delaunay triangulation [38] which is consistent with true facial muscle distribution and uniform for different subjects [3].
 Then each landmark will be connected to the same landmark in the consecutive frame.
@@
 An identity matrix $I$, of the same dimensions as $A$, is created to represent self-connections.
-A spatial-temporal graph, as the basic element of an ST-GCN layer, is implemented as follows [kipf2016semi, yan2018spatial].
+A spatial-temporal graph, as the basic element of an ST-GCN layer, is implemented as follows [21].
 $f_{\text{out}}=\left(\Lambda^{-\frac{1}{2}}((A+I)\odot M)\Lambda^{-\frac{1}{2}}f_{\text{in}}W_{\text{spatial}}\right)W_{\text{temporal}}$ (1)
 where $\Lambda^{ii}=\sum_{j}(A^{ij}+I^{ij})$.
-$M$ is a learnable weight matrix that enables scaling the contributions of a node’s feature to its neighboring nodes [yan2018spatial].
+$M$ is a learnable weight matrix that enables scaling the contributions of a node’s feature to its neighboring nodes [21].
 The input feature map, denoted $f_{\text{in}}$, is the raw coordinates of facial landmarks for the first layer of ST-GCN, and it represents the outputs from previous layers in subsequent layers of ST-GCN.
@@
 This convolution is a standard 2D convolution with a kernel size of $1\times\Gamma$, where $\Gamma$ signifies the temporal kernel size.
-Following an adequate number of ST-GCN layers that perform spatial-temporal graph convolutions as outlined above, the resulting tensor undergoes 2D average pooling, and the final output of the network is generated by a final 2D convolution.
+Following an adequate number of ST-GCN layers, specifically three in this work, that perform spatial-temporal graph convolutions as outlined above, the resulting tensor undergoes 2D average pooling.
+The final output of the network is generated by a final 2D convolution.
 This convolution employs a kernel size of $1\times 1$ and features an output channel dimensionality equal to the number of classes $K$, i.e., the number of engagement levels to be measured.
+An explanation of the detailed architecture of the ST-GCNs for specific datasets can be found in subsection 4.2.
 3.3 Ordinal Engagement classification through Transfer Learning
-The model described above, with a final layer comprising $K$ output channels, tackles the engagement measurement problem as a categorical $K$-class classification problem without taking into account the ordinal nature of the engagement variable [yannakakis2018ordinal, whitehill2014faces, khan2022inconsistencies].
+The model described above, with a final layer comprising $K$ output channels, tackles the engagement measurement problem as a categorical $K$-class classification problem without taking into account the ordinal nature of the engagement variable [39, 40, 10].
 The model could harness the ordinality of the engagement variable to enhance its inferences.
-Drawing inspiration from [frank2001simple, abedi2023affect], a novel ordinal learning framework based on transfer learning is introduced as follows.
+Drawing inspiration from [41, 13], a novel ordinal learning framework based on transfer learning is introduced as follows.
 Training phase– The original $K$-level ordinal labels, $y=0,1,\ldots,K-1$, in the training set are converted into $K-1$ binary labels $y_{i}$ as follows: if $y>i$, then $y_{i}=1$; otherwise, $y_{i}=0$, for $i=0,1,\ldots,K-2$.
@@
 The tensor obtained from this process is then input into $K-1$ pre-trained final 2D convolution layers, each yielding a probability estimate for the test sample being in the binary class $y_{t}=y_{i}$, where $i=0,1,\ldots,K-2$.
-Subsequently, these $K-1$ binary probability estimates are transformed into a single multi-class probability of the sample belonging to class $y=0,1,\ldots,K-1$, as follows [frank2001simple].
+Subsequently, these $K-1$ binary probability estimates are transformed into a single multi-class probability of the sample belonging to class $y=0,1,\ldots,K-1$, as follows [41].
 $p(y_{t}=k)=\begin{cases}1-p(y_{t}\geq 0),&\text{if }k=0,\\ p(y_{t}>k-1)-p(y_{t}\geq k),&\text{if }0<k<K-1,\\ p(y_{t}>K-2),&\text{if }k=K-1.\end{cases}$ (2)
@@
 It reports and discusses the results of multi-class and binary classification of engagement across two datasets.
-4.1 Evaluation Metrics
 Based on the engagement measurement problem at hand, several evaluation metrics are employed.
 In the context of multi-class engagement level classification, metrics such as accuracy and confusion matrix are reported.
-For binary engagement classification, accuracy, Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC), and Area Under the Curve of the Precision and Recall curve (AUC-PR) are utilized.
-Table 1: The distribution of samples in not engaged, barely engaged, engaged, and highly engaged classes in the train and validation sets in the EngageNet dataset [engagenet].
-Engagement Level Train Validation
-Not engaged 1550 132
-Barely engaged 1035 97
-Engaged 1658 273
-Highly Engaged 3740 569
-Total 7983 1071
-Table 2: The distribution of samples in engaged and not engaged classes in the train, validation, and test sets in the IIITB Online SE dataset [thomas2022automatic].
-Engagement Level Train Validation Test
-Not engaged 570 580 570
-Engaged 2620 1080 720
-Total 3190 1660 1290
+For binary engagement classification, accuracy, the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC), and the Area Under the Curve of the Precision and Recall curve (AUC-PR) are utilized.
+In addition, the number of parameters of the models used, memory consumption, and inference time of the proposed method are compared to those of previous methods.
 4.1 Datasets
-Experiments on two video-based engagement measurement datasets were conducted, each presenting unique challenges that further enabled the validation of the proposed method.
-EngageNet The EngageNet dataset [engagenet], recognized as the largest publicly available dataset for video-based student engagement measurement, includes video recordings of 127 subjects participating in virtual learning courses.
+Experiments on two large video-based engagement measurement datasets were conducted, each presenting unique challenges that further enabled the validation of the proposed method.
+EngageNet: The EngageNet dataset [12], recognized as the largest dataset for student engagement measurement, includes video recordings of 127 subjects participating in virtual learning sessions.
 Each video sample has a duration of 10 seconds, with a frame rate of 30 fps and a resolution of $1280\times 720$ pixels.
-Three independent annotators annotated the subjects’ video recordings as four ordinal levels of engagement: not-engaged, barely engaged, engaged, and highly engaged.
-The inter-rater reliability among the annotators was high, with values of 0.73, 0.77, and 0.79, indicating a substantial agreement level.
-The dataset was divided into 7983, 1071, and 2257 samples for training, validation, and testing, respectively, using a subject-independent data split approach [engagenet].
+The subjects’ video recordings were annotated as four ordinal levels of engagement: Not-Engaged, Barely-Engaged, Engaged, and Highly-Engaged.
+The dataset was divided into 7983, 1071, and 2257 samples for training, validation, and testing, respectively, using a subject-independent data split approach [12].
 However, only the training and validation sets were made available by the dataset creators and were utilized for the training and validation of predictive models in the experiments presented in this paper.
-Table 1 details the distribution of the four engagement classes within the training and validation sets.
-Online SE: The Online SE dataset [thomas2022automatic], comprises videos of six students participating in online courses via the Zoom platform.
+The distribution of samples in the four aforementioned classes of engagement in the training and validation sets are 1550, 1035, 1658, and 3740, and 132, 97, 273, and 569, respectively.
+Online SE: The Online SE dataset [42], comprises videos of six students participating in online courses via the Zoom platform.
 These recordings span 10 seconds each, with a frame rate of 24 fps and a resolution of $220\times 155$ pixels.
-Two independent annotators annotated the video recordings as engaged and not engaged using the Human Expert Labeling Process (HELP) [aslan2017human, abedi2023bag].
-The binary engagement labels achieved an inter-rater reliability score of 0.62, denoting a significant level of agreement.
-Following the exclusion of flawed videos [abedi2023bag], the dataset was segmented into 3190, 1660, and 1290 samples for training, validation, and testing, respectively.
-The distribution of engaged and not-engaged samples across the training, validation, and test sets is outlined in Table 2.
+The videos were annotated as either Not-Engaged or Engaged.
+The dataset was segmented into 3190, 1660, and 1290 samples for training, validation, and testing, respectively.
+The distribution of samples in the two aforementioned classes of engagement in the training, validation, and test sets is 570 and 2620, 580 and 1080, and 570 and 720, respectively.
 4.2 Experimental Setting
 The sole information extracted from video frames is facial landmarks, which are analyzed by ST-GCN to determine the engagement level of the student in the video.
-Drawing inspiration from the pioneering works on body-joints-based action analysis [zheng2023skeleton, yan2018spatial], the proposed ST-GCN for facial-landmarks-based engagement measurement is structured as follows.
+Drawing inspiration from the pioneering works on body-joints-based action analysis [21, 43], the proposed ST-GCN for facial-landmarks-based engagement measurement is structured as follows.
 The input facial landmarks are first processed through a batch normalization layer, followed by three consecutive ST-GCN layers with 64, 128, and 256 output channels, respectively.
@@
 The Sigmoid function is also employed for the individual binary classifiers within the ordinal learning framework detailed in subsection 3.3.
-The models are trained using the Adam optimizer with mini-batches of size $16$ and a learning rate of $0.001$ for $300$ epochs.
+The models are trained using the Adam optimizer with mini-batches of size $16$ and an initial learning rate of $0.001$ for $300$ epochs.
 The learning rate is decayed by a factor of $0.1$ every $100$ epochs.
-Table 3: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet]: comparison of various feature sets and classification models from [engagenet] against two configurations of the proposed method - facial landmarks analyzed by ST-GCN and ordinal ST-GCN.
+Table 1: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [12]: comparison of state-of-the-art end-to-end methods and feature-based methods with various feature sets and classification models against two configurations of the proposed method - facial landmarks analyzed by ST-GCN and ordinal ST-GCN.
 Bolded values denote the best results.
-Features Model Accuracy
-Gaze LSTM 0.6125
-Head Pose LSTM 0.6760
-AU LSTM 0.6303
-Gaze + Head Pose LSTM 0.6769
-Gaze + Head Pose + AU LSTM 0.6704
-Gaze CNN-LSTM 0.6060
-Head Pose CNN-LSTM 0.6732
-AU CNN-LSTM 0.6172
-Gaze + AU CNN-LSTM 0.6275
-Head Pose + AU CNN-LSTM 0.6751
-Gaze + Head Pose + AU CNN-LSTM 0.6751
-Gaze TCN 0.6256
-Head Pose TCN 0.6611
-AU TCN 0.6293
-Gaze + Head Pose + AU TCN 0.6779
-Gaze Transformer 0.5545
-Gaze + Head Pose Transformer 0.6445
-Gaze + Head Pose + AU Transformer 0.6910
-Gaze + Head Pose + AU + MARLIN Transformer 0.6849
-Facial Landmarks ST-GCN 0.6937
-Facial Landmarks Ordinal ST-GCN 0.7124
-the proposed method
+Ref. Features Model Accuracy
+[27] End to End Model ResNet + TCN 0.5472
+[30] End to End Model EfficientNet + LSTM 0.5757
+[30] End to End Model EfficientNet + Bi-LSTM 0.5894
+[12] Gaze LSTM 0.6125
+[12] Head Pose LSTM 0.6760
+[12] AU LSTM 0.6303
+[12] Gaze + Head Pose LSTM 0.6769
+[12] Gaze + Head Pose + AU LSTM 0.6704
+[12] Gaze CNN-LSTM 0.6060
+[12] Head Pose CNN-LSTM 0.6732
+[12] AU CNN-LSTM 0.6172
+[12] Gaze + AU CNN-LSTM 0.6275
+[12] Head Pose + AU CNN-LSTM 0.6751
+[12] Gaze + Head Pose + AU CNN-LSTM 0.6751
+[12] Gaze TCN 0.6256
+[12] Head Pose TCN 0.6611
+[12] AU TCN 0.6293
+[12] Gaze + Head Pose + AU TCN 0.6779
+[12] Gaze Transformer 0.5545
+[12] Gaze + Head Pose Transformer 0.6445
+[12] Gaze + Head Pose + AU Transformer 0.6910
+[12] Gaze + Head Pose + AU + MARLIN Transformer 0.6849
+[34] Gaze TCCT-Net 0.6433
+[34] Head Pose TCCT-Net 0.6891
+[34] AU TCCT-Net 0.6629
+[34] Gaze + Head Pose TCCT-Net 0.6564
+[34] Gaze + Head Pose + AU TCCT-Net 0.6713
+Ours Facial Landmarks ST-GCN 0.6937
+Ours Facial Landmarks Ordinal ST-GCN 0.7124
 4.3 Experimental Results
-Table 3 presents the comparative results of two settings of the proposed method, a regular non-ordinal classifier and an ordinal classifier, with previous methods on the validation set of the EngageNet dataset [engagenet].
-The engagement measurement in EngageNet [engagenet] is a four-class classification problem and the accuracy is reported as the evaluation metric.
-The previous methods used different combinations of OpenFace’s eye gaze, head pose, and AU features [baltrusaitis2018openface] along with MARLIN’s facial embedding features [cai2023marlin].
-These features were classified by LSTM, CNN-LSTM, TCN, and Transformer, refer to [engagenet] for more details of the previous methods.
+4.3.1 Comparison to Previous Methods
+Table 1 presents the comparative results of two settings of the proposed method, a regular non-ordinal classifier and an ordinal classifier, with previous methods on the validation set of the EngageNet dataset [12].
+The engagement measurement in EngageNet [12] is a four-class classification problem and the accuracy is reported as the evaluation metric.
+The previous methods in Table 1 include state-of-the-art end-to-end methods, including the combination of ResNet-50 with TCN [27] and the combination of EfficientNet B7 with LSTM and bidirectional LSTM [30], followed by state-of-the-art feature-based methods.
+The previous feature-based methods in Table 1 used different combinations of OpenFace’s eye gaze, head pose, and AU features [31] along with MARLIN’s facial embedding features [11].
+These features were classified by LSTM, CNN-LSTM, TCN, and Transformer.
+Refer to [12] for more details.
+The results of the feature-based method proposed by Vedernikov et al.
+[34] are also reported, where the aforementioned features are classified using a Transformer-based neural network called the Tensor-Convolution and Convolution-Transformer Network (TCCT-Net).
+Despite the abundant data samples in the EngageNet dataset [12] available for training complex neural networks such as ResNet + TCN [27] and EfficientNet B7 + bidirectional LSTM [30], their performance is inferior to that of feature-based methods.
+This highlights the necessity of extracting hand-crafted features or facial landmarks from videos and building classifiers on top of them.
 In the single feature configurations of previous methods in Table 1, head pose achieves better results compared to AUs, which is better than eye gaze.
-While head pose and eye gaze are indicators of behavioral engagement [abedi2023affect, ocumpaugh2015baker], AUs, which are associated with facial expressions and affect [toisoul2021estimation], are indicators of affective engagement.
-Combining these three features is always beneficial since engagement is a multi-component variable that can be measured by affective and behavioral indicators when the only available data is video [abedi2023affect].
-Among classification models, utilizing more advanced models improves accuracy; the Transformer is better than TCN, which is better than CNN-LSTM and LSTM; however, it comes at the cost of increased computational complexity and the number of model parameters.
-It should also be noted that the feature extraction step for the three feature sets requires three complex deep learning models in OpenFace [baltrusaitis2018openface].
-Considering only the classification models, the number of parameters in the Transformer, according to the architecture described in [engagenet], is 104,963,076, while the numbers of parameters in the proposed non-ordinal and ordinal ST-GCN are 861,688, and 861,431, respectively, indicating the efficiency of the proposed method.
-While being lightweight, the proposed method with ordinal STGCN improves the state-of-the-art by 3.1% compared to eye gaze, head pose, and AUs with Transformer.
-Tables 4(a) and 4(b) depict the confusion matrices of the two configurations of the proposed method in Table 3, facial landmarks classified by non-ordinal and ordinal ST-GCN, respectively.
-As shown in Tables 4(a) and 4(b), incorporating ordinality significantly increases the number of correctly classified samples in the first three classes and results in a 2.7% improvement in accuracy compared to its non-ordinal counterpart.
-Table 4: Confusion matrices of the proposed method using facial landmarks with (a) non-ordinal and (b) ordinal ST-GCN on the validation set of the EngageNet dataset [engagenet].
+While head pose and eye gaze are indicators of behavioral engagement [13, 7], AUs, which are associated with facial expressions and affect [19], are indicators of affective engagement.
+Combining these three features is always beneficial since engagement is a multi-component variable that can be measured by affective and behavioral indicators when the only available data is video [13].
+Among classification models, utilizing more advanced models improves accuracy; the Transformer is better than TCN, which is better than CNN-LSTM and LSTM; however, it comes at the cost of increased computational complexity.
+For single features, the Transfomer-based TCCT-Net [34] outperforms other classifiers; however, for multiple feature sets, the vanilla Transformer [12] outperforms the others.
+The proposed ST-GCN in Table 1, which relies solely on facial landmarks without requiring raw facial videos or multiple hand-crafted features, outperforms previous methods.
+Moreover, making the proposed method ordinal further improves the state-of-the-art by 3.1% compared to eye gaze, head pose, AUs, and MARLIN features [11] with the Transformer [12].
+Tables 2(a) and 2(b) depict the confusion matrices of the ordinal and non-ordinal configurations of the proposed method in Table 1.
+As shown, incorporating ordinality significantly increases the number of correctly classified samples in the first three classes and results in a 2.7% improvement in accuracy compared to its non-ordinal counterpart.
+Table 2: Confusion matrices of the proposed method with (a) non-ordinal and (b) ordinal ST-GCN on the validation set of the EngageNet dataset [12].
+(a) Non-ordinal ST-GCN
 Class 1 2 3 4
@@
 4 5 4 45 515
+(b) Ordinal ST-GCN
 Class 1 2 3 4
@@
 4 9 5 44 511
-Table 5 displays the results of different variants of the proposed method on the validation set of the EngageNet [engagenet] dataset.
-In the first two variants, the $x$, $y$, and $z$ coordinates of facial landmarks are converted into multivariate time series and analyzed by an LSTM and TCN.
-The LSTM includes four unidirectional layers with 256 neurons in hidden units and is followed by a $256\times 4$ fully connected layer.
-The parameters of the TCN are as follows: the number of layers, number of filters, kernel size, and dropout rate are 8, 64, 8, and 0.05, respectively.
-While their results are acceptable and better than most of the earlier methods in Table 3, they cannot outperform ST-GCN, the last two columns of Table 3.
-The fact that the accuracy of the LSTM and TCN in Table 5 is higher than those in Table 3 signifies the efficiency of facial landmarks for engagement measurement.
-In the third variant, the $z$ coordinates of facial landmarks are disregarded, and the decrease in accuracy indicates the importance of the $z$ coordinates for engagement measurement.
-Temporal kernel sizes other than 9 in the fourth and fifth variants have a negative impact on the results.
-When engagement measurement is performed using every 2, 4, 8, and 16 frames, instead of every frame, there is a slight decrease in accuracy.
-However, this is a trade-off between accuracy and computation since reducing the frame rate corresponds to a reduction in computation.
-Table 5: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet] for different variants of the proposed method.
+Table 3: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [12] for different variants of the proposed method.
 Variant Accuracy
@@
 Every 16 frames are used 0.6841
-Table 6 presents the results of the proposed method in comparison to previous methods on the test set of the Online SE dataset [thomas2022automatic].
-The earlier methods listed in Table 6 are feature-based, extracting affective and behavioral features from video frames and performing binary classification of the features using TCN in [thomas2018predicting], TCN in [abedi2023affect], LSTM with attention in [chen2019faceengage], and BoW in [abedi2023bag].
-Given that engagement measurement in the Online SE dataset [thomas2022automatic] is posed as a binary classification problem, implementing the ordinal version of the proposed method is not required.
+Hand landmarks are added 0.6956
+The feature extraction step in [34] and [12] involves running multiple deep-learning models in OpenFace [31] to capture eye gaze, head pose, and AUs, as well as another complex network for MARLIN feature embeddings [11].
+While these feature extraction networks are complex for real-time use, the proposed method relies on facial landmarks extracted using the real-time MediaPipe [14, 20].
+Considering only the classification models, the number of parameters in EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], the proposed non-ordinal ST-GCN, and ordinal ST-GCN are 82,681,812, 24,639,236, 1,063,108, 861,688, and 861,431, respectively.
+The memory consumption for EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], and the proposed ordinal ST-GCN are 2268.78, 790.35, 178.00, and 180.96 megabytes, respectively.
+The inference time for classifying a data sample using EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], and the proposed ordinal ST-GCN are 348, 51, 10, and 0.8 milliseconds, respectively.
+This indicates the efficiency of the proposed method, which, while being lightweight and fast, also improves the state-of-the-art.
+4.3.2 Variants of the Proposed Method
+Table 3 displays the results of different variants of the proposed method on the validation set of the EngageNet [12] dataset.
+In the first two variants, the $x$, $y$, and $z$ coordinates of facial landmarks are converted into multivariate time series and analyzed by an LSTM and TCN.
+The LSTM includes four unidirectional layers with 256 neurons in hidden units and is followed by a $256\times 4$ fully connected layer.
+The parameters of the TCN are as follows: the number of layers, number of filters, kernel size, and dropout rate are 8, 64, 8, and 0.05, respectively.
+While their results are acceptable and better than most of the earlier methods in Table 1, they cannot outperform ST-GCN, the last two rows of Table 1.
+The fact that the accuracy of the LSTM and TCN in Table 3 is higher than those in Table 1 signifies the efficiency of facial landmarks for engagement measurement.
+In the third variant, the $z$ coordinates of facial landmarks are disregarded, and the decrease in the accuracy of the non-ordinal ST-GCN indicates the importance of the $z$ coordinates for engagement measurement.
+Temporal kernel sizes other than 9 in the fourth and fifth variants have a negative impact on the results of the non-ordinal ST-GCN.
+When engagement measurement is performed using every 2, 4, 8, and 16 frames, instead of every frame, there is a slight decrease in the accuracy of the non-ordinal ST-GCN.
+However, this is a trade-off between accuracy and computation since reducing the frame rate corresponds to a reduction in computation.
+The last row of Table 3 shows the results of the ordinal ST-GCN when 21 hand landmarks extracted using MediaPipe [14] were added to the facial landmarks.
+The lower accuracy compared to using only facial landmarks is due to two factors.
+Firstly, there is 80% missingness in the hand landmarks due to the absence of hands in the videos.
+Secondly, it indicates that facial landmarks alone are sufficient for engagement measurement, capturing both behavioral and emotional indicators of engagement.
+4.3.3 Results on the Online SE Dataset
+Table 4 presents the results of the proposed method in comparison to previous methods on the test set of the Online SE dataset [42].
+The earlier methods listed in Table 4 are feature-based, extracting affective and behavioral features from video frames and performing binary classification of the features using TCN in [33], TCN in [13], LSTM with attention in [44], and BoW in [32].
+Given that engagement measurement in the Online SE dataset [42] is posed as a binary classification problem, implementing the ordinal version of the proposed method is not required.
 The proposed method, employing facial landmarks with ST-GCN, attains the highest accuracy and AUC PR.
-Table 6: Binary classification accuracy, AUC ROC, and AUC PR of engagement on the test set of the Online SE dataset [thomas2022automatic]: comparison of previous methods against the proposed method.
+Table 4: Binary classification accuracy, AUC ROC, and AUC PR of engagement on the test set of the Online SE dataset [42]: comparison of previous methods against the proposed method.
 Bolded values denote the best results.
 Method Accuracy AUC ROC AUC PR
-[thomas2018predicting] 0.7803 0.8764 0.8008
-[abedi2023affect] 0.7637 0.8710 0.7980
-[chen2019faceengage] 0.7475 0.8890 0.7973
-[abedi2023bag] 0.8191 0.8926 0.9018
+[33] 0.7803 0.8764 0.8008
+[13] 0.7637 0.8710 0.7980
+[44] 0.7475 0.8890 0.7973
+[32] 0.8191 0.8926 0.9018
 Proposed 0.8315 0.8806 0.9131
-5 Conclusion and Future Works
+4.3.4 Interpretation of Results
+Fig.
+1 displays the interpretation of engagement measurements taken using the proposed method by applying Grad-CAM [43] to the ordinal ST-GCN trained on the training set of the EngageNet dataset.
+For visualization purposes, three exemplary frames (out of 300) from the beginning, middle, and end of three data samples annotated as Not-Engaged in the validation set of the EngageNet dataset are shown in Fig.
+1 (a)-(c).
+The facial landmarks extracted through MediaPipe are overlaid on the frames, where the color map of the facial landmarks, from blue to red, depicts the class activation map values of the last ST-GCN layer in the trained ST-GCN.
+The facial landmarks associated with the target class of Not-Engaged at certain frames are colored towards red.
+In Fig.
+1 (a), during the beginning and middle of the video, the first two exemplary frames show the student engaged, with lower class activation map values.
+At the end of the video, when the student is not looking at the camera (computer screen) and is looking elsewhere, landmarks on the iris, eye, and jawline show higher values, resulting in the model classifying the sample as Not-Engaged.
+In Fig.
+1 (b), the student is not paying attention and is playing with their phone.
+Jawline, eye, iris, and eyebrow facial landmarks with red colors have higher class activation map values, resulting in the model classifying the sample as Not-Engaged.
+In Fig.
+1 (c), throughout the entire video sample, the head pose of the student is normal and perpendicular to the camera.
+However, the eyes are almost closed, indicating sleepiness and low arousal.
+This is detected by the higher class activation map values on the eye and iris landmarks, corresponding to the intensity of AU number 45, which indicates how closed the eyes are.
+This resulted in the model classifying the sample as Not-Engaged.
+Figure 1: Interpretation of engagement measurements taken using the proposed method by applying Gradient-weighted Class Activation Mapping (Grad-CAM) to the Spatial-Temporal Graph Convolutional Network (ST-GCN).
+Please refer to the last paragraph of subsection 4.3.4 for further details.
+5 Discussion
 Our research led to the development of a novel deep-learning framework for student engagement measurement.
-In the proposed framework, sequences of facial landmarks are extracted from consecutive video frames and analyzed by ordinal ST-GCNs to make inferences regarding the engagement level of the subject in the video.
-The successful application of our model to the EngageNet [engagenet] and Online SE [thomas2022automatic] datasets not only confirms its efficacy but also establishes a new standard in engagement level classification accuracy, outperforming existing methods.
-As the sole input information to the developed engagement measurement models, the 3D facial landmarks contain information about head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation], which are the key indicators of behavioral and affective engagement.
-The relatively lightweight ST-GCN and its integration with real-time MediaPipe make the proposed framework capable of being deployed on virtual learning platforms and measuring engagement in real time.
-Despite these advancements, the proposed method is not without limitations.
-The interpretability and explainability of the proposed method could be improved through techniques such as Gradient-based Class Activation Maps [zheng2023skeleton] for ST-GCNs [9746621].
-This would facilitate an understanding of which facial landmarks and specific timestamps contribute to certain levels of engagement.
-To improve the performance of the proposed method, the following direction could be investigated: analyzing facial landmarks with more advanced ST-GCN, which are equipped with attention mechanism [lin2023actionlet], applying augmentation techniques to video data before facial landmark extraction [abedi2023cross] or to facial landmark data to improve the generalizability of the ST-GCNs, and incorporating hand landmarks and body joints extracted from video into the developed facial-landmarks-based framework.
+In the proposed framework, sequences of facial landmarks are extracted from consecutive video frames and analyzed by ordinal ST-GCNs to make inferences regarding the engagement level of the student in the video.
+The successful application of our model to the EngageNet [12] and Online SE [42] datasets not only confirms its efficacy but also establishes a new standard in engagement level classification accuracy, outperforming previous methods.
+As the sole input information to the developed engagement measurement models, the 3D facial landmarks contain information about head pose [17], AUs [18], eye gaze [20], and affect [3, 19], which are the key indicators of behavioral and affective engagement.
+The relatively lightweight and fast ST-GCN and its integration with real-time MediaPipe [14] make the proposed framework capable of being deployed on virtual learning platforms and measuring engagement in real-time.
+The proposed method is privacy-preserving and does not require access to personally identifiable raw video data for engagement measurement.
+In a real-world deployment, the cross-platform MediaPipe solution [14], running on a web, mobile, or desktop application, extracts facial landmarks from video data on users’ local devices.
+These non-identifiable facial landmarks are then transferred to a cloud, where they are analyzed by ST-GCNs to measure engagement.
+The interpretability feature of the proposed method, enabled through Grad-CAM, facilitates understanding which facial landmarks, corresponding to behavioral and affective indicators of engagement, contribute to certain levels of engagement.
+It also helps identify the specific timestamps at which these contributions occur.
+This provides instructors with additional information to take necessary actions and promote student engagement.
+A limitation of the proposed method is its reliance on the quality of facial landmarks detected by MediaPipe.
+In the context of engagement measurement in virtual learning sessions, an occluded or absent face, and consequently non-detected facial landmarks, correspond to lower levels of engagement or disengagement.
+Our developed ST-GCN was able to correctly classify most samples with occluded or absent faces, i.e., no facial landmarks, as Not-Engaged.
+To improve the performance of the proposed method, the following direction could be investigated: analyzing facial landmarks with more advanced ST-GCNs, which are equipped with attention mechanisms and trained through contrastive learning techniques and applying augmentation techniques to video data before facial landmark extraction [45] or to facial landmark data to improve the generalizability of ST-GCNs.
 Acknowledgment The authors express their sincere thanks to the Multimodal Perception Lab at the International Institute of Information Technology, Bangalore, India, for their generosity in providing the Online SE dataset, which was instrumental in the execution of our experiments.
-[…]
+This research was funded by the Natural Sciences and Engineering Research Council of Canada.